Balance the given chemical equation using linear equations (gussis Jordan elimination method).





Na₂CO3 + C + N₂ ---> NaCN + CO

1
Expert's answer
2022-03-07T17:25:05-0500

Let's write the equation in the form

a Na2CO3 + b C + c N2 ---> d NaCN + e CO

where a, b, c, d, and e are unknown numbers.


Equate the left and right sides of the equation:

2 a = d for Na

a + b = d + e for C

3 a = e for O

2 c = d for N


or in matrix form



"\begin{bmatrix}
 2 & 0 & 0 & -1 & 0 \\
 1 & 1 & 0 & -1 & -1 \\
 3 & 0 & 0 & 0 & -1 \\
 0 & 0 & 2 & -1 & 0
\end{bmatrix} \cdot 
\begin{bmatrix}
 a \\
 b \\
 c \\
 d \\
 e
\end{bmatrix} = 
\begin{bmatrix}
 0 \\
 0 \\
 0 \\
 0 \\
 0
\end{bmatrix}"

Then using Gauss-Jordan method

"\begin{bmatrix}
 2 & 0 & 0 & -1 & 0 \\
 1 & 1 & 0 & -1 & -1 \\
 3 & 0 & 0 & 0 & -1 \\
 0 & 0 & 2 & -1 & 0
\end{bmatrix} \rarr
\begin{bmatrix}
 1 & 0 & 0 & -1/2 & 0 \\
 1 & 1 & 0 & -1 & -1 \\
 3 & 0 & 0 & 0 & -1 \\
 0 & 0 & 2 & -1 & 0
\end{bmatrix} \rarr
\begin{bmatrix}
 1 & 0 & 0 & -1/2 & 0 \\
 0 & 1 & 0 & -1/2 & -1 \\
 0 & 0 & 0 & 3/2 & -1 \\
 0 & 0 & 2 & -1 & 0
\end{bmatrix} \rarr"


"\begin{bmatrix}
 1 & 0 & 0 & -1/2 & 0 \\
 0 & 1 & 0 & -1/2 & -1 \\
 0 & 0 & 0 & 3/2 & -1 \\
 0 & 0 & 2 & -1 & 0
\end{bmatrix} \rarr
\begin{bmatrix}
 1 & 0 & 0 & -1/2 & 0 \\
 0 & 1 & 0 & -1/2 & -1 \\
 0 & 0 & 2 & -1 & 0 \\
 0 & 0 & 0 & 3/2 & -1
\end{bmatrix} \rarr
\begin{bmatrix}
 1 & 0 & 0 & -1/2 & 0 \\
 0 & 1 & 0 & -1/2 & -1 \\
 0 & 0 & 1 & -1/2 & 0 \\
 0 & 0 & 0 & 3/2 & -1
\end{bmatrix} \rarr"


"\begin{bmatrix}
 1 & 0 & 0 & -1/2 & 0 \\
 0 & 1 & 0 & -1/2 & -1 \\
 0 & 0 & 1 & -1/2 & 0 \\
 0 & 0 & 0 & 1 & -2/3
\end{bmatrix} \rarr
\begin{bmatrix}
 1 & 0 & 0 & 0 & -1/3 \\
 0 & 1 & 0 & 0 & -4/3 \\
 0 & 0 & 1 & 0 & -1/3 \\
 0 & 0 & 0 & 1 & -2/3
\end{bmatrix}"


Therefore if e = 3 then a = 1, b = 4, c =1, d = 2 and the equation is

Na2CO3 + 4C + N2 ---> 2NaCN + 3CO